OpenStudy (anonymous):

simplify the square root of 500

OpenStudy (zepp):

\(\sqrt{500}\) you can find all the divisors \(\sqrt{2*250}=\sqrt{2*5*50}\) and so on, until you get only prime numbers, then you can try to simplify that down. :)

OpenStudy (anonymous):

It goes like this: \[\sqrt{500}=2*\sqrt{125}=10*\sqrt{5}\] Its a matter of simplifying roots out.